OpenStudy (anonymous):

If V(r)=2/5πr^3 ,, V'(r)= ?

OpenStudy (anonymous):

power rule for this one

OpenStudy (anonymous):

\[V'(r)=3\times \frac{2}{5}\pi r^2=\frac{6}{5}\pi r^2\]